OpenStudy (anonymous):

Simplify 2 ? 52 + 4 ? (-3)3.

OpenStudy (anonymous):

what is "?"

OpenStudy (anonymous):

times

OpenStudy (anonymous):

so the question is "simplify 2(52) + 4(-3)(3)" ?

OpenStudy (anonymous):

no. is 2 time 5 the power of 2+ 4 times (-3)3

OpenStudy (anonymous):

"2(5^2)+4(-3)(3)" ?

OpenStudy (anonymous):

Exponents 1st. so 5^2=25, times 2 = 50, 4(-3) = -12, -12(3) = -36; 50 + (-36) = 50-36 = 14

OpenStudy (anonymous):

ok thank you